OpenStudy (anonymous):

how do I solve (4k+5)(k+1)=0 ? @VeritasVosLiberabit

OpenStudy (helder_edwin):

set each factor equal to zero.

OpenStudy (anonymous):

as @helder_edwin said You need to set the factors \[4k+5=0\] and \[k+1=0\] Now solve each for k. This will give you two values for k

OpenStudy (anonymous):

so i think the answer is: 4k=-5 or k=-1 am i correct or nah?

OpenStudy (anonymous):

yes but simply the first one a little more. You can still divide -5 by 4

OpenStudy (anonymous):

simplify*

OpenStudy (anonymous):

okay. so now it is: k=-5/4?

OpenStudy (anonymous):

Well done, that's it

OpenStudy (anonymous):

so your final values should be k= -5/4, -1

OpenStudy (anonymous):

And this is assuming that the problem wants the zeros to your function since you didn't specify in the question.

OpenStudy (anonymous):

Thanks love! and yes that's it.

OpenStudy (anonymous):

alright, you're welcome

OpenStudy (anonymous):

hey, message me on the regular, I like your help. lol